OpenStudy (anonymous):

Write a parameterization for the curve in the xy-plane. The line traced out from the point (3, −1) at t = 0 to the point (1, 6) at t = 1.

OpenStudy (anonymous):

first find equation for the line: y+1=(7/2)x-21/2 y=(7/2)x-23/2 now let x=t so we'll have y=(7/2)t-23/2 x=t <answer

OpenStudy (anonymous):

i need x= and y=

OpenStudy (anonymous):

x=t and y=(7/2)t-23/2